OpenStudy (anonymous):

wat is the best way 2 clear fractions

OpenStudy (anonymous):

multiply the entire thing by the denom.

OpenStudy (anonymous):

yea and then wats the ne3xt step

jhonyy9 (jhonyy9):

for example 1/3 = 3/7x --- 7x=9

OpenStudy (anonymous):

did u timez the 3*3

OpenStudy (amistre64):

there is no next step; once the fractions are cleared, then they are cleared

OpenStudy (anonymous):

what do the fractions have to do with creating a system for the problem?